OpenStudy (anonymous):

(16^-2)-3/4 ....16 in brackets to the power of -2 outside brackets to the power of -3/4

OpenStudy (hoblos):

(16^-2)^-3/4 = 16^(-2)(-3/4) = 16^(3/2)

OpenStudy (anonymous):

how did u do that?

OpenStudy (hoblos):

(x^a)^b = x^(a*b) this is the formula

OpenStudy (anonymous):

thanks:)

OpenStudy (hoblos):

my pleasure =)